A tank with a capacity of 400 gallons originally contains 100 gallons of water with 25 pounds of salt in solution. Water containing 1/2 pound of salt per gallon is entering at a rate of 4 gallons per minute, and the mixture is allowed to flow out of the tank at a rate of 1 gallon per minute. FIND the amount of salt in the tank at any time prior to the instant when the solution begins to overflow.
Answer #1

Solutions given data : v(t) = 400 gal Vola 100 gal. Q(0) 2 25 Tb tank at time tis v (t) = 100+t da .: Sate in - Xate out : dtdifferential The solution of fine art equation as, Q (t) (100++)?: Su(100++* dt S M 900*23 +C • 6100++)? + Q (t) = (100+t) +